# SMF: Template-free and Rig-free Animation Transfer using Kinetic Codes

### v1

@@ 3.4 Temporally Coherent Motion Prediction
We predict the sequences in chunks of consecutive frames, i.e., the given sequence is split into fixed windows each of size ( in our tests). We set the initial state of the NODE as the Corrective Jacobians required for the first frame as,
To make this operable with ANODE [7], we augment the initial state with extra dimensions. We follow the recommendations of the original work and set the extra dimensions to where denotes the number of extra (augmented) dimensions to be added. We use for all our experiments. This lifts the state space to a higher dimensional space where it can more easily predict the required trajectories. Furthermore, as the ANODE incrementally proceeds, the extra dimensions store valuable information available elsewhere in the sequence. This simplifies the corrective function it has to learn (see evaluation). In other words, temporal message passing helps predict the corrective Jacobians. Equation 8 becomes,
@@
where , as defined previously, are the Jacobians of the first frame; is the attention encoding of the current window of Jacobian predictions from Equation 7 and is the attention encoding of the previous window of corrective predictions. We integrate the local changes over time using Euler’s method to obtain at each time as,
Since is in the augmented space, we do a final linear projection with learnable weights, to map it back to the original unaugmented space. Specifically,
@@
These attention networks encode a window of Jacobians into a single encoding. Fixing the encoding sizes to a constant size enables handling any arbitrary window/sequence length without overflowing memory. The encoders distill the current window of Jacobian predictions and previously predicted window of Corrective Jacobians . We pass the output of the attention networks as conditioning to Equation 10 to integrate and obtain the correctives in Equation 11 in the augmented space, before projecting them in Equation 12 to obtain the final correctives. We add the predicted correctives to the posed Jacobians from Equation 7. Finally, the predicted Jacobians are spatially integrated using a differentiable Poisson solve [27], in the coordinate frame of the first frame, to obtain the predicted shape at frame number .
@@
4 Evaluation
We test SMF along multiple axes: (i) generalization to unseen motion; (ii) animation transfer to different characters (shapes and topology; meshes and scans); (iii) motion specification using 3D keypoints, using monocular videos, or sampling/interpolating in the Kinetic Code (latent) space. No rigs or templates were used for any training or tests.
@@ 4 Evaluation
| NJF [2] | 8.94 | 1.07 | 0.42 | 15.95 | 9.41 | 1.14 | 0.49 | 17.91 | 9.06 | 0.98 | 0.54 | 15.95 | 9.42 | 1.12 | 0.44 | 17.26 | 8.86 | 1.06 | 0.43 | 15.96 | 9.17 | 1.19 | 0.74 | 24.56 | 9.47 | 1.17 | 0.49 | 19.10 |
| TRJ [26] | 4.42 | 0.69 | 0.34 | 13.20 | 4.06 | 0.73 | 0.39 | 14.48 | 5.49 | 0.65 | 0.34 | 12.96 | 4.17 | 0.72 | 0.34 | 13.74 | 5.40 | 0.76 | 0.35 | 13.05 | 4.72 | 0.80 | 0.45 | 20.47 | 4.20 | 0.69 | 0.35 | 13.14 |
| Skeleton-free transfer [20] | 4.60 | 0.67 | 0.44 | 13.67 | 3.27 | 0.65 | 0.50 | 13.78 | 5.27 | 0.78 | 0.65 | 19.27 | 5.27 | 0.72 | 0.46 | 13.97 | 4.71 | 0.62 | 0.40 | 12.82 | 3.74 | 0.70 | 0.52 | 15.39 | 3.83 | 0.75 | 0.56 | 16.19 |
@@
| NJF [2] | 25.57 | 9.03 | 1.13 | 45.94 | 25.38 | 9.02 | 1.11 | 45.67 | 26.27 | 8.96 | 1.10 | 45.24 | 31.80 | 10.68 | 1.19 | 45.80 | 29.22 | 9.71 | 1.14 | 45.26 |
| TRJ [26] | 15.23 | 5.86 | 0.87 | 35.09 | 16.91 | 6.62 | 0.85 | 33.54 | 14.22 | 5.20 | 0.64 | 24.95 | 15.99 | 5.94 | 0.73 | 30.25 | 15.17 | 5.39 | 0.70 | 28.62 |
| Skeleton-free transfer [20] | 12.41 | 4.61 | 1.08 | 32.28 | 13.72 | 3.81 | 0.83 | 25.51 | 13.24 | 2.79 | 0.70 | 21.14 | 17.44 | 2.75 | 0.67 | 20.48 | 14.35 | 2.60 | 0.69 | 20.85 |
@@
We train our method (and baselines) on a single dataset comprising of human motion categories from the AMASS dataset [22] with each category containing approximately 6-7 motion sequences. We evaluate our method on unseen motion categories from AMASS, sampled motions from Mixamo, animal motion sequences from DeformingThings4D dataset [19], and in-the-wild monocular video recordings. The AMASS dataset utilizes the SMPL body model [21], which enables generation of motion sequences for diverse body-shapes by varying the shape parameter, . Note that ours does not use this SMPL information during training or inference.
To show that our method works on animals, we test our method on motions from the DeformingThings4D dataset [19], which provides animal 4D meshes as deforming sequences.
4.1 Baselines
We compare our method against recent per frame (pose transfer) methods: NJF [2] and Skeleton-Free Pose Transfer [20]. Due to unavailability of pre-processing code for Skeleton-Free Pose Transfer, we use their pretrained models which were trained on substantially more data including AMASS and Mixamo motions as well as stylized characters. Additionally, it requires the full mesh sequence whereas ours only takes in sparse keypoints as input. We also compare ours with animation transfer methods: TRJ [26] (which uses template) and an ablated version of our method without the Kinetic Code () encoding.
4.2 Metrics
@@
4.3 Target Shapes
We evaluate our method on diverse target shapes varied body shapes sampled from SMPL models, human scans from the FAUST dataset [5], characters from the Mixamo library (e.g., skeleton zombie, triceratop, wolf), and in-the-wild meshes from online 3D repositories (e.g., alien, hole-man). As preprocessing, when applicable, we converted non-manifold meshes to manifold ones.
4.4 Qualitative Results
@@
We present video results on various unseen motion and unseen shapes in our supplementary webpage. Our method produces consistently better generalization to unseen motion categories compared to NJF and TRJ, both of which result in shape distortion and erroneous displacements as they struggle to follow the input (unseen) motion as seen in Figure 5. This is particularly highlighted in scenarios with large displacements, e.g., feet and hands are widened or stretched thin. Note that unlike in the original paper [26], where specialized TRJ models were separately trained for each motion type, we retrained TRJ across all the motion types for a fair comparison.
SMF also generalizes to new shapes of varied body types, including non-humans despite being trained only on humans. It preserves the source motion and target shape and the resultant motion is realistic and free from jitters/artifacts. We show comparative results on unseen motion transfer to in-the-wild target meshes Figure 5. For high-genus shapes, such as the mesh with holes (right half of Figure 5), NJF distorts the shape; TRJ fares comparatively better, it is still riddled with artifacts. In contrast, our Kinetic codes not only preserve the target shape but also more faithfully adhere to the source motion.
Utilizing a latent representation for motion encoding with smooth interpolation properties leads to improved generalization. Hence, we do not see any significant artifacts even when operating on in-the-wild unseen target shapes for unseen motion. Moreover, we notice without our motion encoding, regions around joints may show melting (see around feet in Figure 3). The same holds true for more accurate bending of the joints (see Figure 5).
Figure 1 shows more animation transfer examples to different target meshes as well as support for different types of coarse motion specifications. In Figure 4, we show retargetting results on target animal shapes.
@@
To transfer motion from Mixamo [24], we map its joint locations (3D keypoints) to our system, (note the mapping is not perfect due to misalignment between joint locations, requiring some approximation). Our results demonstrate strong generalization, transferring these motions even to different characters. Notably, TRJ [26] and NJF [2] are unable to operate on Mixamo motion, as they strictly depend on the SMPL template, which significantly limits their flexibility. Therefore, we modify their architecture slightly to handle 3D coordinates and remove the module from TRJ. Please see supplemental videos on the webpage for comparison. We further evaluate it on 5 different motions on 3 stylized characters with ground truth generated from Mixamo. As seen in Table 2, we see large improvements over existing methods. All methods except Skeleton-free transfer [20] have been trained solely on AMASS motions and human characters.
@@
can directly transfer raw motion sequences from 2D input videos captured on handheld cameras to a 3D target shape. We use our formulation based on 2D keypoints to directly operate on 2D input frames. This is a particularly challenging problem as we are lifting motion from 2D to a 3D shape without any explicit supervision or template shape. For our 2D training setup, we additionally train only the autoencoder on in-the-wild videos of exercise videos totaling 1 hour. We note this is noisy but this helps the generalize to RGB captures better as all our 2D renderings of AMASS sequences have a similar camera setup. We extract 2D keypoints using HRNet [38], which computes 2D keypoints frame by frame. We use Savitzky-Golay filter to smoothen the 2D signal as there is inter-frame noise.
@@ 4.5 Quantitative Comparison
We present quantitative results on unseen motion category and unseen motion sequence (within the same category) to unseen shape in Tables 1 and 2. Generalization to unseen motion is a particularly challenging problem as the Poisson solve in NJF and TRJ facilitates shape preservation for unseen shape by following the coordinates of the source shape. However, there is no such aid when generalizing to unseen motion, which is an extrapolation problem. Our method significantly outperforms NJF and TRJ. This holds even for long motion sequences (1000-4000+ frames) such as Walk. The benefits of our is highlighted here as we see significantly lower errors against a baseline of without encoding. Compared to the per-frame pose transfer performed by Skeleton-Free Pose Transfer which has numerous artifacts and unnatural deformations, SMF also ensures temporal consistency as seen in lower velocity error (L2-V), measuring motion smoothness.
5 Conclusion
@@
We have presented Self-supervised Motion Fields that convert sparse motion into full-body character motion. We enable this by first creating a temporally-informed motion latent space – Kinetic codes – and then utilizing it to train spatial and temporal gradient predictor networks jointly. The gradient signals are coupled via spatial and temporal integration, and trained using full-body mesh sequences for supervision. SMF does not require additional motion annotation, is simple and robust to train, and generalizes across unseen motion and character shape variations. We also do not use any template model or rig, either during training or inference. We establish a new SOTA on the AMASS dataset by showing improvements over related approaches.
@@
Limitations. While our neural ODE formulation captures some secondary motion, we do not model secondary dynamics in any physically correct way. Neither do we explicitly support collision handling to prevent self-intersection. An interesting future direction would be to implicitly handle collision detection and simultaneously model character-garment interactions, possibly using a transformer-based attention mechanism to capture non-local interactions between body parts.
Future work. We would like to take the output as structure-guidance to condition video generators to produce stylized videos. We believe such a 3D-consistent scaffolding should lead to improved object permanence and deformations in the output videos, and may be key to generate controllable and interpretable outputs.

### v2

@@ 3.4. Temporally Coherent Motion Prediction
We introduce a second stage that learns to predict a dynamic correction to the initial pose estimate to enforce temporal coherence. We predict the sequences in chunks of consecutive frames, i.e., the given sequence is split into fixed windows, each of size ( in our tests). We set the initial state of the NODE as the Corrective Jacobians required for the first frame as,
To make this operable with ANODE (Dupont et al., 2019), we augment the initial state with extra dimensions. We follow the recommendations of the original work and set the extra dimensions to where denotes the number of extra (augmented) dimensions to be added. We use for all our experiments. In essence, these augmented dimensions act as a form of temporal memory for the ODE. By lifting the state space to a higher dimensional space where it can more easily predict the required trajectories, the system is able to maintain a richer state representation of the motion’s history. This simplifies the task of learning the corrective dynamics needed to prevent drift. In other words, temporal message passing helps predict the corrective Jacobians. Equation 8 becomes,
@@
where , as defined previously, are the Jacobians of the first frame; is the attention encoding of the current window of Jacobian predictions from Equation 7 and is the attention encoding of the previous window of corrective predictions. Intuitively, provides the context of the current motion’s structure, while informs the model about the accumulated error from the previous window. We integrate the local changes over time using Euler’s method to obtain at each time as,
Since is in the augmented space, we use a final linear projection with learnable weights () to map it back to the original unaugmented space. Specifically,
@@
These attention networks encode a window of Jacobians into a single encoding as shown in Figure 3. Fixing the encoding sizes to a constant size enables handling any arbitrary window/sequence length without overflowing memory. The encoders distill the current window of Jacobian predictions and previously predicted window of Corrective Jacobians . We pass the output of the attention networks as conditioning to Equation 10 to integrate and obtain the correctives in Equation 11 in the augmented space, before projecting them in Equation 12 to obtain the final correctives. We add the predicted correctives to the posed Jacobians from Equation 7. Finally, the predicted Jacobians are spatially integrated using a differentiable Poisson solve (Nicolet et al., 2021), in the coordinate frame of the first frame, to obtain the predicted shape at frame number .
@@
4. Evaluation
We test SMF along multiple axes: (i) generalization to unseen motion; (ii) animation transfer to different characters (shapes and topology; meshes and scans); (iii) motion specification using 3D keypoints, using monocular videos, or sampling/interpolating in the Kinetic Code (latent) space. No rigs or templates were used for any training or tests. We also test motion transfer from models trained on humanoid characters to non-humanoids.
@@ 4. Evaluation
| Skeleton-free transfer (Liao et al., 2022) | 4.60 | 0.67 | 0.44 | 13.67 | 3.27 | 0.65 | 0.50 | 13.78 | 5.27 | 0.78 | 0.65 | 19.27 | 5.27 | 0.72 | 0.46 | 13.97 | 4.71 | 0.62 | 0.40 | 12.82 | 3.74 | 0.70 | 0.52 | 15.39 | 3.83 | 0.75 | 0.56 | 16.19 |
@@
| Skeleton-free transfer (Liao et al., 2022) | 12.41 | 4.61 | 1.08 | 32.28 | 13.72 | 3.81 | 0.83 | 25.51 | 13.24 | 2.79 | 0.70 | 21.14 | 17.44 | 2.75 | 0.67 | 20.48 | 14.35 | 2.60 | 0.69 | 20.85 |
| (Aigerman et al., 2022) | 25.57 | 9.03 | 1.13 | 45.94 | 25.38 | 9.02 | 1.11 | 45.67 | 26.27 | 8.96 | 1.10 | 45.24 | 31.80 | 10.68 | 1.19 | 45.80 | 29.22 | 9.71 | 1.14 | 45.26 |
| (Muralikrishnan et al., 2024) | 15.23 | 5.86 | 0.87 | 35.09 | 16.91 | 6.62 | 0.85 | 33.54 | 14.22 | 5.20 | 0.64 | 24.95 | 15.99 | 5.94 | 0.73 | 30.25 | 15.17 | 5.39 | 0.70 | 28.62 |
@@
We train our method (and baselines) on a single dataset comprising of human motion categories from the AMASS dataset (Mahmood et al., 2019) with each category containing approximately 6-7 motion sequences. We evaluate our method on unseen motion categories from AMASS, sampled motions from Mixamo, animal motion sequences from DeformingThings4D dataset (Li et al., 2021b), and in-the-wild monocular video recordings. The AMASS dataset utilizes the SMPL body model (Loper et al., 2015), which enables generation of motion sequences for diverse body-shapes by varying the shape parameter, . Note that ours does not use this SMPL information during training or inference.
To show that our method works on animals, we train our method on motions from the DeformingThings4D dataset (Li et al., 2021b), which provides animal 4D meshes as deforming sequences and test it on unseen motion sequences.
We train SMF on 9 motion categories from AMASS, each consisting of sequences performed by 6 humans. The number of frames varies from 150-800. For 2D keypoints, we use 3 YouTube videos (2 characters in total), totaling 1hr. Although the videos are noisy as the camera angles change, SMF successfully learns due to the Kinetic codes setup.
4.1. Baselines
We compare our method against recent per frame (pose transfer) methods: NJF (Aigerman et al., 2022) and Skeleton-Free Pose Transfer (Liao et al., 2022). Due to unavailability of pre-processing code for Skeleton-Free Pose Transfer, we use their pretrained models which were trained on substantially more data including AMASS and Mixamo motions as well as stylized characters. Additionally, it requires the full mesh sequence along with T pose meshes for source and target characters. In comparison, ours only takes in sparse keypoints as input. We also compare ours with animation transfer methods: TRJ (Muralikrishnan et al., 2024) (which uses template) and an ablated version of our method without the Kinetic Code () encoding. In the table below, we highlight the differences between SMF and baseline methods.
| Method | Input Requirements |
| NJF (Aigerman et al., 2022) | SMPL Pose Parameters |
| Skeleton Free (Liao et al., 2022) | T-Posed Source & Target characters, full body source motion |
| TRJ (Muralikrishnan et al., 2024) | SMPL Pose and Bodyshape Parameters |
| SMF (ours) | Sparse Keypoints |
| Modified & trained with Keypoints | |
| Modified & trained with Keypoints |
4.2. Metrics and Target Shapes
@@
Target Shapes
We evaluate our method on diverse target shapes varied body shapes sampled from SMPL models, human scans from the FAUST dataset (Bogo et al., 2014), characters from the Mixamo library (e.g., skeleton zombie, triceratop, wolf), and in-the-wild meshes from online 3D repositories (e.g., alien, hole-man). As preprocessing, when applicable, we fixed non-manifold meshes.
4.3. Qualitative Results
@@
We present video results on various unseen motion and unseen shapes on our supplementary webpage. Our method produces consistently better generalization to unseen motion categories compared to NJF and TRJ, both of which result in shape distortion and erroneous displacements as they struggle to follow the input (unseen) motion as seen in Figure 4. This is particularly highlighted in scenarios with large displacements, e.g., feet and hands are widened or stretched thin. Note that unlike in the original TRJ (Muralikrishnan et al., 2024), where specialized models were separately trained for each motion type, we retrained a single TRJ, across all the motion types.
SMF also generalizes to new shapes of varied body types, including non-humans despite being trained only on humans. This correspondence from coarse keypoints to a dense mapping across diverse shapes is learned during the self-supervised motion transfer setup and proves to be even capable of generalizing to multi-legged creatures. It preserves the source motion and target shape and the resultant motion is realistic and free from jitters/artifacts. We show comparative results on unseen motion transfer to in-the-wild target meshes in Figure 4. For high-genus shapes, such as the mesh with holes (right half of Figure 4), NJF distorts the shape; TRJ fares comparatively better, it is still riddled with artifacts. Our Kinetic codes not only preserve the target shape but also more faithfully adhere to the motion. See supplemental webpage for video results.
Utilizing a latent representation for motion encoding with smooth interpolation properties leads to improved generalization. Hence, we do not see any significant artifacts even when operating on in-the-wild unseen target shapes for unseen motion. Moreover, we notice without our motion encoding, regions around joints may show melting (see around feet in Figure 6). The same holds true for more accurate bending of the joints (see Figure 4). We note that the AMASS motions contain foot-skating artifacts, and ours faithfully reproduces them. However, we also tested ours on high-quality Mixamo motions as shown in Figure 4 and supplemental webpage, where foot skating is less pronounced. Despite being trained only on AMASS, our model successfully transfers these clean Mixamo motions without introducing any additional skating artifacts. This demonstrates the model’s generalization not only to unseen motions but also to different levels of motion quality. This further suggests that our self-supervised setting is promising and training on higher-quality motion samples would likely yield better results.
Figure 1 and Figure 8 show more animation transfer examples to different humanoids and non-humanoids target meshes, sourced from online in-the-wild meshes and the SHREC’07 dataset (Giorgi et al., 2007). Moreover, Figure 1 showcases support for different types of coarse motion specifications. In Figure 7, we show retargetting results on target animal shapes from the D4D dataset.
@@
To transfer motion from Mixamo (Mixamo, 2025), we map its joint locations (3D keypoints) to our system, (note the mapping is not perfect due to misalignment between joint locations, requiring some approximation). Our results demonstrate strong generalization, transferring these motions even to different characters. Notably, TRJ (Muralikrishnan et al., 2024) and NJF (Aigerman et al., 2022) are unable to operate on Mixamo motion, as they strictly depend on the SMPL template, which significantly limits their flexibility. Therefore, we modify their architecture to handle 3D coordinates and remove the shape parameter () module from TRJ. Please see supplemental videos on the webpage for comparison. We further evaluate it on 5 different motions on 3 stylized characters with ground truth generated from Mixamo. As seen in Table 2, we see large improvements over existing methods. Except Skeleton-free transfer (additionally trained on Mixamo data and stylized characters), all other baselines have been trained solely on the AMASS dataset with human characters.
@@
can directly transfer raw motion sequences from 2D input videos captured on handheld cameras to a 3D target shape. We use our formulation based on 2D keypoints to directly operate on 2D input frames. This is a particularly challenging problem as we are lifting motion from 2D to a 3D shape without any explicit supervision or template shape. For our 2D training setup, we additionally train only the autoencoder on in-the-wild videos of exercise videos totaling 1 hour. We note this is noisy but this helps the generalize to RGB captures better as all our 2D renderings of AMASS sequences have a similar camera setup. We extract 2D keypoints using HRNet (Sun et al., 2019), which computes 2D keypoints frame by frame. We use Savitzky-Golay filter to smoothen the 2D signal as there is inter-frame noise.
@@ 4.4. Quantitative Comparison
We present quantitative results on unseen motion category and unseen motion sequence (within the same category) to unseen shape in Tables 1 and 2. Generalization to unseen motion is a particularly challenging problem as the Poisson solve in NJF and TRJ facilitates shape preservation for unseen shape by following the coordinates of the source shape. However, there is no such aid when generalizing to unseen motion, which is an extrapolation problem. Our method significantly outperforms NJF and TRJ for both motion accuracy as well as detail preservation. The benefits of our are highlighted here as we see significantly lower errors against a baseline of without encoding. Compared to the per-frame pose transfer performed by Skeleton-Free Pose Transfer which has numerous artifacts and unnatural deformations, SMF also ensures temporal consistency as seen in lower velocity error (L2-V), measuring motion smoothness.
Error accumulation on long motion transfer.
We choose 3 unseen, complex, and long dance sequences (10000 frames) to analyze error accumulation. We report per frame vertex-to-vertex (L2-V) error in Figure 9. Compared to the only other temporal method, TRJ, SMF does not exhibit a trend of monotonically increasing error demonstrating its robustness to temporal drift and its suitability for generating long motion sequences. The initial spike in error is due to the motion moving away from the easy-to-model rest pose.
4.5. Performance
SMF’s lightweight architecture enables inference at ¿30 FPS on a single RTX 4090 for a mesh with 30k faces (for high number of faces this will be slower). This makes our system suitable for real-time applications, following a one-time preprocessing step to compute features on the target mesh. Furthermore, when using 2D keypoints, the system could potentially be extended for real-time 2D-to-3D animation transfer, as keypoint extraction with HRNet (Sun et al., 2019) is highly efficient.
5. Conclusion
@@
We have presented Self-supervised Motion Fields that convert sparse motion into full-body character motion. We enable this by first creating a temporally-informed motion latent space – Kinetic codes – and then utilizing it to train spatial and temporal gradient predictor networks jointly. The gradient signals are coupled via spatial and temporal integration, and trained using full-body mesh sequences for supervision. SMF does not require additional motion annotation, is simple and robust to train, and generalizes across unseen motion and character shape variations. We also do not use any template model or rig, either during training or inference. We establish a new SOTA on the AMASS dataset.
@@
Limitations. While our formulation captures some secondary motion, we do not model secondary dynamics in any physically correct way. Neither do we explicitly support collision handling to prevent self-intersection (see the triceratop leg crossing in Figure 7). An interesting future direction would be to implicitly handle collision detection and simultaneously model character-garment interactions, possibly using a transformer-based attention mechanism to capture non-local interactions between body parts. Similar to single human bodies (Frühstück et al., 2022), it will be interesting to combine different motion generators for different parts of a character, allowing mixing of motion modes. Finally, temporal sketching sequences (Fu et al., 2011) can also be modeled with such a generative framework.